OpenStudy (anonymous):

Find the surface area of the pyramid to the nearest whole number 8 in length 7.2 slant height

OpenStudy (anonymous):

\(\Large S=\frac{1}{2}\times 8\times \sqrt{7.2^2-4^2} \approx 24 \)